# Highly Efficient Self-Adaptive Reward Shaping for Reinforcement Learning

### v1

@@ 6 Conclusion and Discussion
We propose SASR in this paper, a self-adaptive reward shaping algorithm based on success rates, to address the sparse-reward challenge. SASR achieves an exploration and exploitation trade-off mechanism by generating shaped rewards from evolving Beta distributions. Experiments demonstrate that this adaptability significantly enhances the agent’s convergence speed. Additionally, the implementation of KDE and RFF formulates a highly efficient, non-parametric, and learning-free approach to deriving Beta distributions. This mechanism also provides a sound alternative to traditional count-based RS strategies, adapting effectively to continuous environments. Our evaluations confirm the superior performance of SASR in terms of sample efficiency and learning stability, and its robustness across various continuous-control tasks.
While SASR is designed for environments with sparse rewards, in settings where the environmental rewards are dense, the introduction and maintenance of additional shaped rewards could be unnecessary. On the other hand, SASR operates on the premise that sparse rewards indicate task completion to align the success rate notion. This correlation is generally straightforward, yet it may not intuitively apply in all contexts. Exploring how to extend SASR to more general sparse and even dense reward scenarios presents a promising direction for further research. Moreover, the derivation of Beta distributions heavily relies on the samples in the success and failure buffers. Currently, our method does not capture the relationships and varying importance of different states within the same trajectory, making the algorithm relatively sensitive to the retention rate, a crucial hyperparameter for achieving promising performance. Therefore, developing an adaptive retention rate or better mechanisms for maintaining the two buffers are crucial avenues for future enhancements.
@@ Appendix A Derivation of Random Fourier Features
We incorporate Random Fourier Features (RFF) [Rahimi and Recht, 2007] to approximate the kernel functions in the KDE process for the SASR algorithm. Let the original state be $k$-dimensional, denoted as $\bm{s}\in\mathbb{R}^{k}$, and the kernel function be $k(\bm{s}_{i},\bm{s}_{j})$. RFF approximates the kernel function by projecting the input $k$-dimensional space into a $M$-dimensional feature space using a mapping function $z:\mathbb{R}^{k}\rightarrow\mathbb{R}^{M}$. The RFF-based kernel function is then defined as follows:
@@
Gaussian $\exp\left(-\frac{\|\bm{s}_{i}-\bm{s}_{j}\|_{2}^{2}}{2}\right)$ $(2\pi)^{-\frac{M}{2}}\exp\left(-\frac{\|\bm{w}\|_{2}^{2}}{2}\right)$
@@ Appendix B Derivation of Computation Complexity
Appendix C Supplementary Experimental Results for Ablation Study
@@
Appendix D Network Structures and Hyperparameters
D.1 Network Structures
@@ D.2 Hyperparameters
Appendix E Compute Resources

### v3

@@ 6 Conclusion and Discussion
In this paper, we propose SASR, a self-adaptive reward shaping algorithm based on success rates, to address the sparse-reward challenge. SASR achieves an exploration and exploitation balance mechanism by generating shaped rewards from evolving Beta distributions. Experiments demonstrate that this adaptability significantly enhances the agent’s convergence speed. Additionally, the implementation of KDE and RFF formulates a highly efficient, non-parametric, and learning-free approach to deriving Beta distributions. This mechanism also provides a sound alternative to traditional count-based RS strategies, adapting effectively to continuous environments. Our evaluations confirm the superior performance of SASR in terms of sample efficiency and learning stability.
While SASR is designed for sparse-reward environments, in settings with dense rewards, the introduction and maintenance of additional shaped rewards could be unnecessary. Exploring the extension of SASR to dense reward scenarios presents a promising direction for further research. Moreover, the derivation of Beta distributions depends heavily on the samples stored in the success and failure buffers. Currently, our method does not account for the relationships or varying importance of different states within the same trajectory, making the algorithm sensitive to the retention rate $\phi$. Therefore, developing an adaptive retention rate or better mechanisms for managing the buffers are crucial avenues for future improvement.
7 Acknowledgement
@@ Appendix A Derivation of Random Fourier Features
We incorporate Random Fourier Features (RFF) (Rahimi and Recht,, 2007) to approximate the kernel functions in the KDE process for the SASR algorithm. Let the original state be $k$-dimensional, denoted as $\bm{s}\in\mathbb{R}^{k}$, and the kernel function be $k(\bm{s}_{i},\bm{s}_{j})$. RFF approximates the kernel function by projecting the input $k$-dimensional space into a $M$-dimensional feature space using a mapping function $z:\mathbb{R}^{k}\rightarrow\mathbb{R}^{M}$. The RFF-based kernel function is then defined as follows:
@@
Gaussian $\exp\left(-\frac{\|\bm{s}_{i}-\bm{s}_{j}\|_{2}^{2}}{h^{2}}\right)$ $\frac{\sqrt{h}}{2\sqrt{\pi}}\exp\left(-\frac{h\|\bm{w}\|_{2}^{2}}{4}\right)$
@@ Appendix B Derivation of Computation Complexity
Appendix C Comparison of SASR with and without RFF
To evaluate the effect of introducing RFF, we compare the training time of SASR with and without RFF, also with the backbone SAC algorithm, the results are shown in Table 3. The tests are conducted on the NVIDIA RTX A6000 GPUT. The results show that excluding SAC’s inherent optimization time, RFF significantly saves time in the SASR algorithm, while with varying effects across tasks.
Table 3: Comparison of training time (in hours) for SASR with and without RFF.
Algorithms AntStand AntFar HumanStand HumanKeep RobotReach RobotPush
SAC (backbone) 5.87 5.08 4.87 5.67 5.42 6.3
SASR KDE+RFF 7.15 7.52 6.92 6.20 7.07 8.13
SASR w/o RFF 8.12 8.72 8.37 6.53 11.12 9.21
Appendix D Supplementary Experimental Results for Ablation Study
@@
Appendix E Network Structures and Hyperparameters
E.1 Network Structures
@@ E.2 Hyperparameters
Appendix F Compute Resources
@@
Appendix G Configurations of Tasks
We provide the detailed configurations of the tasks in the experiments in this section.
  • AntStand: The ant robot is trained to stand over a target position. The reward is given if the ant robot reaches the target height. Maximum episode length is 1000 steps.
  • AntSpeed: The ant robot is trained to move as fast as possible. The reward is given if the ant robot reaches a target speed. Maximum episode length is 1000 steps.
  • AntFar: The ant robot is trained to reach a target position far from the starting point. The reward is given if the ant robot reaches the target position. Maximum episode length is 1000 steps.
  • AntVeryFar: The ant robot is trained to reach a target position very far from the starting point. The reward is given if the ant robot reaches the target position. Maximum episode length is 1000 steps.
  • WalkerKeep: The walker robot is trained to keep a target height. The robot is initialized by standing, and the reward is given if the robot maintains the target height. Maximum episode length is 1000 steps.
  • HumanStand: The human robot is trained to stand over a target position. The robot is initialized by lying on the ground, and the reward is given if the robot reaches the target height. Maximum episode length is 1000 steps.
  • HumanKeep: The human robot is trained to keep a target height. The robot is initialized by standing, and the reward is given if the robot maintains the target height. Maximum episode length is 1000 steps.
  • RobotReach: The robot arm is trained to reach a target position. The target position is randomly generated in the workspace, and the reward is given if the robot reaches the target position. Maximum episode length is 500 steps.
  • RobotPush: The robot arm is trained to push an object to a target position. The target position is randomly generated on the table, and the reward is given if the object reaches the target position. Maximum episode length is 500 steps.
  • MountainCar: The car is trained to reach the top of the right hill. The reward is given if the car reaches the top. Maximum episode length is 1000 steps.